# Grover Speedup from Many Forms of the Zeno Effect

### v3

@@ 6 Numerical Methods
All numerical calculations were performed using Python, making use of numpy [53] for simple matrix operations, and scipy[54] (in particular scipy.linalg.eigh) for the numerical matrix diagonalisation required for figure 10 (a symmetric subspace was used so the Hamiltonian matrix was of size $21$x21212121 and therefore did not require sparse diagonalisation methods). For all single avoided crossing calculations, the exact expressions for eigenvectors and eigenvalues were used, and therefore numerical diagonalisation was not required. Matplotlib[55] was used extensively for plotting.
@@ 7 Discussion
We now step back for a moment to summarise what we have shown. In each of the three families of Zeno effects we have studied, we have found three manifestations of the effect which can all yield optimal speedup for unstructured search. These are summarized in table 1, including the underlying equations describing the dynamics we examine in this work. By virtue of the nature of the scaling arguments, this also implies that interpolations of the from given in [36] will be possible for the continuous versions of both the decoherence and destruction families of algorithms.
@@
Table 1: Equations generating manifestations of different families of Zeno effects. In this work we have shown that all manifestations given here are capable of yielding an optimal Grover speedup on unstructured search. To fit the table within the page width we have used some shorthand $U(t)=\exp(-ig(\tau)/2\tilde{Z})$, $\Pi_{de}=|d\rangle\!\langle e|$, $\Pi_{ed}=\Pi^{\dagger}_{de}=|e\rangle\!\langle d|$. In the leftmost column fd is shorthand for “full discrete”, pd for “partial discrete”, and cont for “continuous time”. We have also shortened “continuous time quantum walk” to ctqw. As a reminder of terms we have defined elsewhere in the paper $\tilde{Z}=|g\rangle\!\langle g|-|e\rangle\!\langle e|$, $\Pi_{e}=|e\rangle\!\langle e|$, $\Pi_{g}=|g\rangle\!\langle g|$. Note that the first two entries for the destruction family do not conserve probability, it is implied that the “missing” probability is the probability to be in the computationally useless state $|d\rangle$.
@@
Comparing figures 2, 6, and 8 in the discrete case and 3, 7, and 9 in the continuous case, we observer a pattern in the constant factors on the scaling for our very simplified model. We find that the phase rotation family has the most favourable constant factors, followed by decoherence, followed by destruction. However, two remarks should be made here, firstly it is not clear whether or not these factors would translate to structured (optimisation) problems. Secondly, even if they did, if it were easier to implement decoherence or destruction manifestations of a Zeno effect (see for example [12]), then it could still be preferable to apply those algorithms in a real setting.
@@ 8 Conclusion
First and foremost we have demonstrated that within a physically realistic setting, quantum Zeno effects due to decoherence (or equivalently measurement) and decay into a computationally useless state, are able to support an optimal quantum speedup for unstructured search. This is an important step because it demonstrates that analog optimisers, in the spirit of quantum annealers, but using these effects to achieve a Zeno effect are well motivated. This is not a given since these quantum channels are often considered undesirable and to detract from the quantum nature of a device. Discussing how such a device could actually be achieved is beyond the scope of this work, as our current focus is on what is theoretically achievable.
It is important to note that this extends beyond previous work which has shown that for example universal gate-model quantum computation can be based on Zeno-effect gates [5]. In these settings the effects would typically be limited to one or two qubit subspaces. What we have shown is that analog quantum computation where a many-body manifestations of the effects are present can also support a quantum advantage. Such specialized devices are likely to be easier to build in the near term, in analogy to the setting of quantum annealing.
@@ 9 Acknowledgements
All authors were fully supported by Quantum Computing Inc. in completing this work. The authors thank Milan Begliarbekov and Yong Meng Sua for useful discussions and references and Uchenna Chukwu for spotting some small errors in the text.
@@
Appendix A Unstructured Search Hamiltonian
@@
A further extension is to qudits, where each variable has dimension $d$. Assume a driver which has equal hopping between adjacent states $|i\rangle\leftrightarrow|i+1\rangle$. The graph formed by this driver within the solution space is again a hypercube, but includes interior points, not just corners. In the context of quantum walks, it has been shown that a Grover speedup is possible for hypercubes of dimension four and greater ($n\geq 4$ in our notation), with an optimal speedup for $n>4$ [33] and $d\geq 2$. This structure for quantum walk implies a description with a single avoided crossing model which will be discussed in the next section, so our result extends to qudits as well, except for the case of three or fewer qudits. Computing with a fixed number of qudits does not represent a valid way to perform scalable quantum computing [61], so these cases are not interesting in understanding computational advantage.
Appendix B Single Avoided Crossing Model
Although it is mathematically harder to show [33, 36], the relevant search dynamics when using a transverse field driver as given in Eq. 46, become increasingly well described as a two level system, with an avoided crossing having a gap proportional to $N^{-1/2}=2^{-\frac{n}{2}}$, where the two-level system implies that $N=2^{n}$. As discussed in the last section, this also extends to qudits for $n>4$, with a gap proportional to ${N}^{-1/2}=d^{-\frac{n}{2}}$, where $N=d^{n}$ in this case. For this reason we are able to consider a model which captures multiple cases, the single avoided crossing model from [36], which was shown in that paper to be valid for a transverse field driver as the size becomes asymptotically large, and by virtue of the two-level nature of the fully connected graph driver Hamiltonian, is also a valid description of that Hamiltonian in the large size limit. A similar approach, but without an explicit model was also used in [49]. Additionally this structure is present in more general hypercubes which describe the solution space of qudit systems [33]. In the limit of large systems, this model can be expressed in the terms of the previously defined Pauli $X$ operators and the Pauli $Z$, $Z=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$. The approach to this limiting behaviour has been extensively studied in [36] The generally time dependent Hamiltonian in the single avoided crossing approximation takes the form
@@ Appendix B Single Avoided Crossing Model
where $T$ is the total runtime and $g_{\mathrm{min}}$ is the minimum energy gap between the ground and first excited state. Since the schedule is defined in terms of the runtime as a fraction of the total rather than the absolute runtime we will henceforth define a rescaling $t/T\rightarrow\tau$ and use $\tau$ as the relevant time parameter. The physical intuition behind this Hamiltonian, as explained in [36] is that one can effectively “scale to infinity” all behaviour except for directly at the avoided crossing. For this reason the boundaries are moved to $f(0)=\infty$ and $f(1)=-\infty$. An advantage of this model is that it applies equally well to the ground and first excited state of any model where the evolution becomes dominated by a single avoided crossing. This applies to both drivers we have discussed here, as well as much more widely to many other cases [49]. As an example of the approach to single avoided crossing behaviour, figure 10 shows the spectrum of the (qubit) hypercube Hamiltonian for $n=20$, a size which already strongly shows the limiting behaviour. A minor disadvantage of the single avoided crossing model is that it loses information about constant factors away from the asymptotic limit (as were examined in detail in [36]). Since the purpose of our current work is to argue the existence of speedups, not to compare models, this is not a major drawback.
Figure 10: Approach to single avoided crossing behaviour for a hypercube search with $n=20$. The y-axis depicts the energy of the eigenstates, while the x-axis depicts the parameter $s$ for A Hamiltonian of the form $H=(1-s)H_{\mathrm{tf}}+ns|m\rangle\!\langle m|$. Here $H_{\mathrm{tf}}$ comes from equation 46. Insets are zooms of the same plot with the right box showing a zoom within the box depicted on the main figure and the lower inset showing a zoom on the right plot. Color indicates eigenstate probability to be measured in relevant states, with red indicating the marked state and blue indicating $|\tilde{\omega}\rangle$; yellow indicates any other state.
@@
where $g(\tau)$ is the energy gap between the ground and first excited state defined in equation 48. We have mapped the initial state $|\tilde{\omega}\rangle\rightarrow|1\rangle$ in this model and $|m\rangle\rightarrow|0\rangle$. Applying the boundary conditions and solving we obtain,
@@
Appendix C Operating principles of Zeno and non-Zeno algorithms
C.1 Purity arguments for measurements and decoherence
In the case we are considering, in which the density matrices are two dimensional, the density matrix can always be written as

### v4

@@ 6 Numerical Methods
All numerical calculations were performed using Python, making use of numpy [54] for simple matrix operations, and scipy[55] (in particular scipy.linalg.eigh) for the numerical matrix diagonalisation required for figure 10 (a symmetric subspace was used, so the Hamiltonian matrix was of size $21$x21212121 and therefore did not require sparse diagonalisation methods). For all single avoided crossing calculations, the exact expressions for eigenvectors and eigenvalues were used, and therefore numerical diagonalisation was not required. Matplotlib[56] was used extensively for plotting.
@@ 7 Discussion
We now step back for a moment to summarise what we have shown. In each of the three families of Zeno effects we have studied, we have found three manifestations of the effect, which can all yield optimal speedup for unstructured search. These are summarised in table 1, including the underlying equations describing the dynamics we examine in this work. By virtue of the nature of the scaling arguments, this also implies that interpolations of the form given in [37] will be possible for the continuous versions of both the decoherence and destruction families of algorithms.
@@
Table 1: Equations generating manifestations of different families of Zeno effects. In this work, we have shown that all manifestations given here are capable of yielding an optimal Grover speedup on unstructured search. To fit the table within the page width we have used some shorthand $U(t)=\exp(-ig(\tau)/2\tilde{Z})$, $\Pi_{de}=|d\rangle\!\langle e|$, $\Pi_{ed}=\Pi^{\dagger}_{de}=|e\rangle\!\langle d|$. In the leftmost column, fd is shorthand for “full discrete”, pd for “partial discrete”, and cont for “continuous time”. We have also shortened “continuous time quantum walk” to ctqw. As a reminder of terms we have defined elsewhere in the paper $\tilde{Z}=|g\rangle\!\langle g|-|e\rangle\!\langle e|$, $\Pi_{e}=|e\rangle\!\langle e|$, $\Pi_{g}=|g\rangle\!\langle g|$. Note that the first two entries for the destruction family do not conserve probability, it is implied that the “missing” probability is the probability to be in the computationally useless state $|d\rangle$.
@@
Comparing figures 2, 6, and 8 in the discrete case and 3, 7, and 9 in the continuous case, we observe a pattern in the constant factors on the scaling for our very simplified model. We find that the phase rotation family has the most favourable constant factors, followed by decoherence, followed by destruction. However, two remarks should be made here, firstly it is not clear whether or not these factors would translate to structured (optimisation) problems. Secondly, even if they did, if it were easier to implement decoherence or destruction manifestations of a Zeno effect (see for example [12]), then it could still be preferable to apply those algorithms in a real setting.
@@
While the results we have shown are interesting from a theoretical perspective, a question we have not fully addressed thus far is what practical relevance the models we have discussed here have. In particular, while theoretically useful, all the Hamiltonians we have discussed here would be more complicated to implement than directly implementing an unstructured search Hamiltonian (for example using the methods of [61]). A natural question is whether there is actually a physical situation where it may be easier to implement either a dephasing or destruction model. We discuss at least one technology which answers this question in the affirmative for the dissipation model.
In section 1 we previewed the potential relevance of this work to the entropy computing paradigm through quantum Zeno blockade. The realisation of a Zeno blockade involves an optical cavity (for example a micro-ring resonator) which contains or is made of a non-linear optical material which supports sum-frequency generation [62]. There also has to be a freely propagating mode which starts outside the cavity (for example, an optical fibre coupled to the cavity). The Hamiltonian of the simplest model of the Zeno blockade (ignoring the photon propagation in the fibre as a simplification for illustrative purposes) is therefore [^4] Here we use standard quantum optics notation, where all operators are given hats, including Hamiltonians:
$\hat{H}_{\mathrm{ZB}}=c\left(\hat{a}_{f}\hat{a}^{\dagger}_{\mathrm{cav}}+\hat{a}^{\dagger}_{f}\hat{a}_{\mathrm{cav}}\right)+G\left(\hat{a}_{cav}\hat{a}_{\mathrm{cont}}\hat{a}^{\dagger}_{\mathrm{loss}}+\hat{a}_{\mathrm{cav}}^{\dagger}\hat{a}^{\dagger}_{cont}\hat{a}_{\mathrm{loss}}\right)$ (41)
From an engineering perspective, it is difficult to make a cavity which is highly resonant with three modes at once [63, 64]. As a result, it is natural to also consider loss from one of the modes (the mode corresponding to the operator $\hat{a}_{\mathrm{loss}}$). This results in a total master equation of the form
$\frac{\partial\hat{\rho}}{\partial t}=-i[\hat{H}_{\mathrm{ZB}},\hat{\rho}]+\gamma\left[\hat{a}_{\mathrm{loss}}\hat{\rho}\hat{a}^{\dagger}_{\mathrm{loss}}-\frac{1}{2}(\hat{a}^{\dagger}_{\mathrm{loss}}\hat{a}_{\mathrm{loss}}\hat{\rho}+\hat{\rho}\hat{a}^{\dagger}_{\mathrm{loss}}\hat{a}_{\mathrm{loss}})\right].$ (42)
In this setting, the Zeno effect is modulated by the off diagonal density matrix element which captures the coherence between a freely propagating mode (corresponding to $\hat{a}_{f}$) and the captured cavity mode (corresponding to $\hat{a}_{\mathrm{cav}}$). In fact, only the imaginary component of this term contributes to transfer of a freely propagating photon into the cavity, this leads to (assuming $m$ photons are present in the control mode and $n$ in the freely propagating mode),
$\frac{\partial}{\partial t}\mathrm{Im}\left[\rho_{cf}\right]=-\sqrt{n}c\left(\rho_{ff}-\rho_{cc}\right)-\sqrt{m\,n}\,G\,\mathrm{Re}\left[\rho_{lf}\right].$ (43)
For the density matrix element $\mathrm{Re}\left[\rho_{lf}\right]$, we have the formula
$\frac{\partial}{\partial t}\mathrm{Re}\left[\rho_{lf}\right]=\sqrt{m\,n}\,G\,\mathrm{Im}\left[\rho_{cf}\right]+\sqrt{n}\,c\,\mathrm{Im}\left[\rho_{lc}\right]-\frac{\gamma}{2}\mathrm{Re}\left[\rho_{lf}\right].$ (44)
If we assume that an effective Zeno blockade is implemented, then terms proportional to $c$ can be neglected, leading to an overall combined differential equation of
$\frac{\partial^{2}}{\partial t^{2}}\mathrm{Im}\left[\rho_{cf}\right]=-m\,n\,G^{2}\mathrm{Im}\left[\rho_{cf}\right]-\frac{\gamma}{2}\frac{\partial}{\partial t}\mathrm{Im}\left[\rho_{cf}\right],$ (45)
which is the same underlying differential equation as a damped harmonic oscillator. The term proportional to the loss rate $\gamma$ acts as an effective damping term, When $4\,\sqrt{n\,m}\,G>\gamma$, this equation is effectively under damped and $\mathrm{Im}\left[\rho_{cf}\right]$ will oscillate between positive and negative values, as it decays, characteristic in a phase rotation manifestation of the Zeno effect. On the other hand if $4\,\sqrt{n\,m}\,G\leq\gamma$ the system is over (or critically) damped and $\mathrm{Im}\left[\rho_{cf}\right]$, then this element will decay but not change sign, as would occur in a destruction based manifestation of the Zeno effect. These regimes roughly map to the coherent (CQZ) and incoherent (IQZ) quantum Zeno blockade regimes as discussed in previous literature [13, 62, 65] This regime is also the one which is more likely to be realised with small $m$ and $n$, in particular in the $m=n=1$ case of a single-photon Zeno blockade. This few or single photon regime is the one which is most relevant to entropy computing [12].
@@ 8 Conclusion
First and foremost we have demonstrated that within a physically realistic setting, quantum Zeno effects due to decoherence (or equivalently measurement) and decay into a computationally useless state, are able to support an optimal quantum speedup for unstructured search. This is an important step because it demonstrates that analog optimisers, in the spirit of quantum annealers, but using these effects to achieve a Zeno effect are well motivated. This is not a given, since these quantum channels are often considered undesirable and to detract from the quantum nature of a device. Discussing how such a device could actually be achieved is beyond the scope of this work, as our current focus is on what is theoretically achievable.
It is important to note that this extends beyond previous work which has shown that for example universal gate-model quantum computation can be based on Zeno-effect gates [5]. In these settings, the effects would typically be limited to one or two qubit subspaces. What we have shown is that analog quantum computation where a many-body manifestations of the effects are present can also support a quantum advantage. Such specialized devices are likely to be easier to build in the near term, in analogy to the setting of quantum annealing.
@@ 9 Acknowledgements
All authors were fully supported by Quantum Computing Inc. in completing this work. The authors thank Milan Begliarbekov and Yong Meng Sua for useful discussions and references, and Uchenna Chukwu for spotting some small errors in the text.
@@
Appendix A Review of the model from [47]
Since the work in our present manuscript builds on a model originally presented in [47] it is worth reviewing the key aspects of the model of measurement from that paper, to make our paper more self-contained. It is worth noting that [47] briefly mentioned Zeno based computation by measurement early in the paper, but does not analyse the potential for a quantum speedup. In fact, the protocol they discuss is one based on linear interpolation, which would not yield a quantum speedup for search [33].
The model we are interested in was rather used to motivate the amount of time over which a single measurement could be performed. This time is shown to scale as $\sqrt{N}$. The measurement time combined with the fact that after this measurement there is an $O(1)$ success probability independent of system size, implies that an algorithm based on single measurements is theoretically able to achieve the same scaling as Grover’s famous gate model algorithm [31, 32]. A key fact to note is that the model of measurement from [47] is not necessarily meant to replicate a real physical device, so much as to show that it is physically realistic to be able to measure on a timescale proportional to $\sqrt{N}$. Our argument follow a similar line of reasoning, to explore what is possible, rather than propose a concrete design. It is worth noting that the Hamiltonian used in this model can be viewed as a continuous-time variant of gate model phase estimation [66, 67]. This connection was made explicit in section 4 of [68].
The key idea behind this model is that an arbitrary von Neumann measurement can be represented as time evolution under a Hamiltonian which couples as system to a continuous variable, which is then measured to obtain the value. Let us assume the goal is to measure the expectation value of a Hermitian operator $\hat{A}$. We will follow the notation of [68] [^5] While we do not use hats for operators elsewhere, it is useful to use them here for clarity and consistency with other literature, we do however retain the dimensionless units of time and energy, which effectively remove factors of $\hbar$. and consider measuring an eigenstate of $\hat{A}$, $|\Psi_{a}\rangle$ with an eigenvalue of $a$. This leads to the property that
$\hat{A}|\Psi_{a}\rangle=a|\Psi_{a}\rangle.$ (46)
We further define a continuous system initialised at a single position, $|x\rangle$. As usual this system will have position and momentum operators $\hat{X}$ and $\hat{P}$ respectively, and these will obey the commutation relationship $\left[\hat{X},\hat{P}\right]=i$. A measurement can then be conceived of as effectively evolving with a Hamiltonian of the form
$\hat{H}=-k\hat{A}\hat{P}.$ (47)
Applying time evolution with respect to this Hamiltonian yields
$\hat{U}(t)|\Psi_{a}\rangle|x\rangle=|\Psi_{a}\rangle|x+k\,a\,t\rangle.$ (48)
By the linearity of quantum mechanics, applying this operator to a superposition of such eigenstates will lead to a superposition of position values.
However, we must recall a key property of measurement is continuous systems is that the position cannot be measured with infinite precision (both because $x$ eigenstates cannot be prepared perfectly, and because of limitations of the apparatus, we consider the former in the main text for convenience). This precision is clearly independent of any properties of $\hat{A}$, or the runtime of the experiment. If we assume a fixed precision at which $x$ measurements can be reliably performed $\Delta x$, and our aim is to differentiate states with eigenvalues $a_{0}<a_{1}$, then the required time for the measurement will be
$t=\frac{\Delta x}{k(a_{1}-a_{0})},$ (49)
in other words, it will depend inversely on the difference in expectation values of $\hat{A}$, which is the energy gap in the case where $\hat{A}$ is a Hamiltonian.
It is worth briefly noting that the evolution described here can be simulated in a gate model setting. In this setting, a convenient way to implement the time evolution operator $\hat{U}(t)$ is in the momentum basis rather than the real space basis [66, 67]. In this implementation, a register is initialised in $|x=0\rangle$ which in the momentum basis is an equal positive superposition of all bitstrings. Each bit is then used as a control on a small angle unitary of with rotation angles proportional to different binary numbers. This imprints phases on the binary-encoded momentum state. An inverse quantum Fourier transform then converts these into a binary representation of $x$ which can be measured. This implementation is the famed phase estimation algorithm, it is reviewed for example in section 5 of [67], but we will not discuss the details here since this work focuses on the continuous-time setting.
Appendix B Unstructured Search Hamiltonian
@@
A further extension is to qudits, where each variable has dimension $d$. Assume a driver which has equal hopping between adjacent states $|i\rangle\leftrightarrow|i+1\rangle$. The graph formed by this driver within the solution space is again a hypercube, but includes interior points, not just corners. In the context of quantum walks, it has been shown that a Grover speedup is possible for hypercubes of dimension four and greater ($n\geq 4$ in our notation), with an optimal speedup for $n>4$ [34] and $d\geq 2$. This structure for quantum walk implies a description with a single avoided crossing model which will be discussed in the next section, so our result extends to qudits as well, except for the case of three or fewer qudits. Computing with a fixed number of qudits does not represent a valid way to perform scalable quantum computing [69], so these cases are not interesting for understanding computational advantage.
Appendix C Single Avoided Crossing Model
Although it is mathematically harder to show [34, 37], the relevant search dynamics when using a transverse field driver as given in Eq. 55, become increasingly well described as a two level system, with an avoided crossing having a gap proportional to $N^{-1/2}=2^{-\frac{n}{2}}$, where the two-level system implies that $N=2^{n}$. As discussed in the last section, this also extends to qudits for $n>4$, with a gap proportional to ${N}^{-1/2}=d^{-\frac{n}{2}}$, where $N=d^{n}$ in this case. For this reason we are able to consider a model which captures multiple cases, the single avoided crossing model from [37], which was shown in that paper to be valid for a transverse field driver as the size becomes asymptotically large, and by virtue of the two-level nature of the fully connected graph driver Hamiltonian, is also a valid description of that Hamiltonian in the large size limit. A similar approach, but without an explicit model, was also used in [50]. Additionally, this structure is present in more general hypercubes which describe the solution space of qudit systems [34]. In the limit of large systems, this model can be expressed in the terms of the previously defined Pauli $X$ operators and the Pauli $Z$, $Z=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$. The approach to this limiting behaviour has been extensively studied in [37] The generally time dependent Hamiltonian in the single avoided crossing approximation takes the form
@@ Appendix C Single Avoided Crossing Model
where $T$ is the total runtime and $g_{\mathrm{min}}$ is the minimum energy gap between the ground and first excited state. Since the schedule is defined in terms of the runtime as a fraction of the total rather than the absolute runtime, we define a rescaling $t/T\rightarrow\tau$ and use $\tau$ as the relevant time parameter.
The physical intuition behind this Hamiltonian, as explained in [37] is that one can effectively “scale to infinity” all behaviour except for directly at the avoided crossing. For this reason, the boundaries are moved to $f(0)=\infty$ and $f(1)=-\infty$. An advantage of this model is that it applies equally well to the ground and first excited state of any setting where the evolution becomes dominated by a single avoided crossing. This applies to both drivers we have discussed here, and much more widely to many other cases [50]. As an example of the approach to single avoided crossing behaviour, figure 10 shows the spectrum of the (qubit) hypercube Hamiltonian for $n=20$, a size which already strongly shows the limiting behaviour. A minor disadvantage of the single avoided crossing model is that it loses information about constant factors away from the asymptotic limit (as were examined in detail in [37]). Since the purpose of our current work is to argue the existence of speedups, not to compare models, this is not a major drawback.
Figure 10: Approach to single avoided crossing behaviour for a hypercube search with $n=20$. The y-axis depicts the energy of the eigenstates, while the x-axis depicts the parameter $s$ for A Hamiltonian of the form $H=(1-s)H_{\mathrm{tf}}+ns|m\rangle\!\langle m|$. Here $H_{\mathrm{tf}}$ comes from equation 55. Insets are zooms of the same plot, with the right box showing a zoom within the box depicted on the main figure and the lower inset showing a zoom on the right plot. Color indicates eigenstate probability to be measured in relevant states, with red indicating the marked state and blue indicating $|\tilde{\omega}\rangle$; yellow indicates any other state.
@@
where $g(\tau)$ is the energy gap between the ground and first excited state defined in equation 57. We have mapped the initial state $|\tilde{\omega}\rangle\rightarrow|1\rangle$ in this model and $|m\rangle\rightarrow|0\rangle$. Applying the boundary conditions and solving, we obtain,
@@
Appendix D Operating principles of Zeno and non-Zeno algorithms
D.1 Purity arguments for measurements and decoherence
In the case we are considering, in which the density matrices are two-dimensional, the density matrix can always be written as